OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (−5, −5) and a directrix of y = 7

OpenStudy (jdoe0001):

have you covered parabolas yet?

hero (hero):

Insert the points (-5 ,-5) and (x , 7) into this formula \((x - x_1)^2+ (y - y_1)^2 = (x - x_2)^2 + (y - y_2)^2\) \((x + 5)^2 + (y + 5)^2 = (y - 7)^2\) Now Expand: \(x^2 + 10x + 25 + y^2 + 10y + 25 = y^2 -14y + 49\) \((x + 5)^2 -24 = -10y - 14y\) \((x + 5)^2 - 24 = -24y\) \(-\dfrac{(x + 5)^2 - 24}{24} = y\) \(y = -\dfrac{(x + 5)^2}{24} + 1\)

OpenStudy (anonymous):

oh ok @Hero